OpenStudy (anonymous):

A lunch special involves selling only fish, medium portions of chips, and potato cakes over a luncheon periods from 12 noon to 2:30pm. The portions were priced as follows: Flake: $6.00 Chips: $4.50 Potato cake: $1 On the first occasion the lunch special was trialled, 255 portions of food were sold. The total takings for the luncheon special were $1067.50. It was known that there were 50 more portions of chips sold than there were pieces of flake. a) Construct a system of linear simultaneous equations to represent this.

OpenStudy (anonymous):

I know two of them will be: f+c+p = 255 and 6f+4.5c+p = 1067.50 but im having trouble with the third one?

OpenStudy (kainui):

Well it says there were 50 more portions of chips sold than pieces of potato cake (I'm assuming that's what pieces of flake is). So c=50+p is your last equation. You can rearrange it to looks like this: 0f+c-p=50